OpenStudy (anonymous):

The density of ice is 0.92g/cm3nd power. An ice sculptor orders a 1.0-m3nd power block of ice. what is the mass of the block? Hint: 1m3=1,000,000 cm3. Give your answer in grams and kilograms.

OpenStudy (happinessbreaksbones):

@hartnn @experimentX @thomaster @ganeshie8 maybe they will help you.

OpenStudy (shane_b):

\[\frac{0.92g}{1\cancel{cm^3}}*\frac{1,000,000\cancel{cm^3}}{1m^3}=\frac{920,000g}{1m^3}=\frac{920kg}{1m^3}\]So the mass of the block is 920,000 grams...or 920kg.